Four consecutive multiples of 10 have a sum of 300. What is the least of these numbers ?
(A) 20 (B) 30 (C) 40 (D) 50 (E) 60

Answer by vksarvepalli(154)

let the no.s be 10n, 10(n+1), 10(n+2), 10(n+3)
given their sum is 300
so 10(n+n+1+n+2+n+3)=300
=> 10(4n+6)=300
=> 4n+6=30
=> 4n=24
so n=6
so the least number is 10n i.e., 60
ans: E